On January 1, 2018, a corporation issued a $100,000, 5-year maturity, 10% bond. The interest is...

  1. On January 1, 2018, a corporation issued a $100,000, 5-year maturity, 10% bond. The interest is payable semi- annually on June 30 and December 31. The market interest rate is 10% at issuance. What is the interest expense for the year ending December 31, 2018?
    A. $12,000

    B. $10,741

    C. $10,000

    D. $6,000

  2. E. $5,386

plz use computer and plz explain with good calculations.. plz
Answer #1

C. $10,000

When coupon rate and market rate is same, price of bond is equal to its par value.So, coupon interest paid for the year is equal to interest expense for the year.

Step-1:Calculation of Issued proceeds from bond
Issued proceeds =-PV(rate,nper,pmt,fv) Where,
= $ 1,00,000.00 rate = 5%
nper = 10
pmt = $             5,000
fv = $       1,00,000
Step-2:Calculation of interest expense for year ended 2018
Interest for first half = 100000*5% = $       5,000
Interest for second half = 100000*5% = $       5,000
Total $     10,000